2 litre of oxygen at 0.7 bar and 0.5 litre of nitrogen at 0.8 bar are put together in a flask of 1 litre capacity at the same temperature . What is the final pressure of the mixture in the flask

Grade:11

1 Answers

Arun
5 years ago
Use gaseous formula , 
PV = nRT => n = PV/RT
now, 
moles of N2 = PV/RT 
for N2, pressure ( P ) = 0.8 bar and volume ( V ) = 0.5 L ,
moles of N2 = 0.8 × 0.5/RT = 0.4/RT
similarly for O2, 
pressure ( P ) = 0.7 bar , volume ( V ) = 2 L
moles of O2 = PV/RT 
= 0.7 × 2/RT = 1.4/RT 
hence, total number of moles = 0.4/RT + 1.4/RT 
= 1.8/RT 
now, Total pressure = n'RT/V' [ by gas formula]
here, n' is Total moles = 1.8/RT 
V' is total volume = 1 L
now, total pressure = 1.8/RT × RT/1 = 1.8 atm